# Multi-Agent Guided Policy Search for Non-Cooperative Dynamic Games

### v3

@@ II Related works
Model-based solutions to non-cooperative games. Computing Nash equilibria of non-cooperative dynamic games is NP-hard [14]; however, under smoothness assumptions on dynamics and costs, they can be solved in polynomial time [1]. For instance, LQ games enable efficient computation, with feedback Nash equilibria obtained via coupled Riccati equations or KKT conditions [8]. Extending these solutions to general nonlinear games remains difficult. Recent advances include iLQGames, which applies local LQ models and dynamic programming in an iterative scheme [7], and differential dynamic programming [6]. Other approaches compute equilibria using Newton’s method for KKT systems [15, 8] or approximate dynamic programming [16, 17], but high-dimensional nonlinear games remain challenging.
Data-driven MARL methods. MARL has been extensively studied in cooperative settings [18, 19, 20, 21, 11], with notable successes in actor–critic methods [22, 23, 24, 25], value decomposition [26, 27], and graph neural network approaches [28]. In contrast, learning game-theoretic equilibria [29] (e.g., Nash equilibria) with deep RL remains challenging due to coupled dynamics and conflicting objectives. Policy gradient methods converge reliably only in special classes such as two-agent zero-sum games [30, 31] and potential games [32], where agents’ costs are symmetric. In general non-cooperative games, however, they may converge to limit cycles rather than a Nash equilibrium [5]. Furthermore, directly applying guided policy search [33], which improves training in single-agent RL, to multi-agent Nash equilibrium settings is nontrivial due to misaligned objectives, except in fully cooperative cases [10, 11].
@@ IV Analysis: Stability and bias of the guided policy optimization in LQ games
We denote $\hat{K}^{*}=[\hat{K}^{1*},\dots,\hat{K}^{N*}]$ as the Nash equilibrium solution for this guiding-policy-regularized LQ game. We show that the introduced regularization term can resolve the instability of the MA-PG dynamics (7).
@@
These observations suggest that $\rho$ trades off stability and bias. If the guidance matches the true Nash equilibrium, additional policy gradient updates become unnecessary; otherwise, the converged solution under the guided policy gradient dynamics can outperform an imperfect guidance policy $\check{K}$.
@@ V-A Efficient computation of local guidance for MA-PG
Unlike the LQ scenario, where a stabilizing policy can be easily computed as guidance, it is generally hard to obtain guidance for MA-PG for nonlinear games due to the complexity of dynamics. Inspired by the iLQGames method [7], which argues that along the Nash equilibrium state and control trajectories of a nonlinear game, the Nash equilibrium of the locally constructed LQ game along those state and control trajectories can approximate the Nash equilibrium of the original nonlinear game well. Building upon this idea, we construct local LQ game approximations along the neural network control output and the resulted state trajectory, and then use the Nash equilibrium of those local LQ game approximations as reference signal to guide the learning of multi-agent policy gradients.
@@ VI Experiments
We evaluate our method against MA-PG methods in three examples: 1) a two-agent linear quadratic (LQ) game, 2) a platooning scenario involving three vehicles merging into a single lane, and 3) a 24-dimensional six-player basketball formation, where agents strategically move to desired positions based on their roles. Implementation details [^2] Code is available at https://github.com/qugechen/MA-GPS of experiments are included in Appendix.
@@
As shown in Figure 3, MA-GPS noticeably increases the convergence rate and reduces variance compared to other four methods. As observed in the trajectory plot, as the number of agents increases, it becomes increasingly difficult for all agents to explore and find an effective policy together. Our guidance provides an effective prior for all agents, helping them converge to a reasonable equilibrium.

### v4

@@ II Related works
Model-based solutions to non-cooperative games. Computing Nash equilibria of non-cooperative games is computationally challenging [7]; however, under smoothness assumptions on dynamics and costs, they can be solved in polynomial time [3]. For instance, LQ games enable efficient computation, with feedback Nash equilibria obtained via coupled Riccati equations or KKT conditions [20]. Extending these solutions to general nonlinear games remains difficult. Recent advances include iLQGames, which applies local LQ models and dynamic programming in an iterative scheme [12], and differential dynamic programming [9]. Other approaches compute equilibria using Newton’s method for KKT systems [11, 20] or approximate dynamic programming [10, 30], but high-dimensional nonlinear games remain challenging.
Data-driven MARL methods. MARL has been extensively studied in cooperative settings [38, 5, 37, 29, 22], with notable successes in actor–critic methods [17, 1, 8, 40], value decomposition [35, 31], and graph neural network approaches [28]. In contrast, learning game-theoretic equilibria [24] (e.g., Nash equilibria) with deep RL remains challenging due to coupled dynamics and conflicting objectives. Policy gradient methods and their variants converge reliably in two-agent zero-sum games [42, 6, 34] and potential games [15], where agents’ costs are symmetric. In general non-cooperative games, however, they may converge to limit cycles rather than a Nash equilibrium [26]. Furthermore, directly applying guided policy search [21], which improves training in single-agent RL, to multi-agent Nash equilibrium settings is nontrivial due to misaligned objectives, except in fully cooperative cases [16, 22].
@@ IV Analysis: Stability and bias of the guided policy optimization in LQ games
We denote $\hat{K}^{*}=[\hat{K}^{1*},\dots,\hat{K}^{N*}]$ as the Nash equilibrium solution for this guiding-policy-regularized LQ game. We show that the introduced regularization term can resolve the instability of the MA-PG dynamics (7), thereby guaranteeing local exponential convergence to a stationary point of the regularized game.
@@
These observations suggest that $\rho$ trades off stability and bias. Moreover, if the guidance matches the true Nash equilibrium, additional policy gradient updates become unnecessary; otherwise, the converged solution under the guided policy gradient dynamics can outperform an imperfect guidance policy $\check{K}$.
@@ V-A Efficient computation of local guidance for MA-PG
Unlike the LQ scenario, where a stabilizing policy can be easily computed as guidance, it is generally hard to obtain guidance for MA-PG for nonlinear games due to the complexity of dynamics. Inspired by the iLQGames method [12], which argues that along the Nash equilibrium state and control trajectories of a nonlinear game, the Nash equilibrium of the locally constructed LQ game along those state and control trajectories can approximate the Nash equilibrium of the original nonlinear game well. Building upon this idea, we construct local LQ game approximations along the neural network control output and the resulting state trajectory, and then use the Nash equilibrium of those local LQ game approximations as reference signal to guide the learning of multi-agent policy gradients.
@@ VI Experiments
We evaluate our method against MA-PG methods in three examples: 1) a two-agent linear quadratic (LQ) game, 2) a platooning scenario involving three vehicles merging into a single lane, and 3) a 24-dimensional six-player basketball formation, where agents strategically move to desired positions based on their roles. Although the last two examples fall outside Assumption 1, our method remains effective, suggesting that the stability analysis of Theorem 1 may extend beyond the LQ setting. Implementation details [^2] Code is available at https://github.com/qugechen/MA-GPS of experiments are included in Appendix.
@@
As shown in Figure 3, MA-GPS noticeably increases the convergence rate and reduces variance compared to other four methods, particularly during the early stages of training, highlighting the stabilizing effect of the proposed model-based guidance. As observed in the trajectory plot, as the number of agents increases, it becomes increasingly difficult for all agents to explore and find an effective policy together. Our guidance provides an effective prior for all agents, helping them converge to a reasonable equilibrium.